Suppose a random sample of n = 16 observations is selected from a population that is normally distributed with mean equal to 101 and standard deviation equal to 12.

QUESTION: Find the probability that the sample mean deviates from the population mean μ = 101 by no more than 4. (Round your answer to four decimal places.)
Answer #1

Sell If x= N(402) then z=x-u Noir Given: mean full =101 standard deviation (0)=12 sample size (n)-16 sample mean (Mist) - Ma